### 22 Jan 2024 v2

# Semi-supervised learning for White Blood Cell Semantic Segmentation

Abstract
AI’s application in healthcare, especially in white blood cell cancer diagnosis, is hindered by two primary challenges: the lack of large-scale labeled datasets for white blood cell (WBC) segmentation and outdated segmentation methods. To address the first challenge, we must have a semi-supervised learning framework so that we can efficiently annotate that large dataset. In this work, we address this issue and propose a novel self-training pipeline with the incorporation of FixMatch. We conduct experiments on white blood cell datasets and evaluate the strengths and weaknesses of our pipeline.
Keywords:
Semi-supervised learning Semantic segmentation White blood cell
1 Introduction
Cancer stands among the deadliest diseases in the world. In Vietnam and the U.S., cancer is only second to heart disease in terms of mortality rate. In 2020, it is estimated that nearly 19 million cases of cancer arise worldwide, with the number of deaths standing at approximately 10 million [1]. Cancer of the white blood cells (WBC), namely leukemia and lymphoma, are two of the most common types of blood cancers [2]. However, certain screening and diagnosis of blood cancers, particularly cancer of the white blood cells (WBC), still employ a manual process, such as in peripheral blood smear analysis. These tests often serve the purpose of classifying and counting the number of WBCs in the blood, and if the number is abnormally high or low, further examination is needed to decide if the patient is inflicted with cancer or not.
White blood cell segmentation is a rarely touched area among AI experts. In our experience, it is hard to find a dataset that both provides images and their ground truth masks to perform semantic segmentation. Most datasets concern themselves with white blood cell classification and do not contain the masks needed for segmentation. And when we found datasets that suited our purposes, the number of data was very insignificant.
Recognizing that there is a shortage of meaningful datasets in the field of white blood cell segmentation, we aim to provide a semi-supervised learning technique so that in the presence of a large-scale dataset, we can annotate that dataset most efficiently.
Our contributions to this paper are as follows:
    We aim to provide a semi-supervised learning approach to annotate a large dataset of white blood cells.
    From that semi-supervised learning approach, we should devise a more accurate, fine-tuned way to segment white blood cells.
2 Related works
2.0.1 Semi-supervised learning.
is a paradigm in which the dataset used for training is not fully annotated. It sits between supervised learning, where the whole dataset is labeled, and unsupervised learning, where no data is labeled. In the early days, semi-supervised learning was confined to traditional machine-learning approaches [5], such as semi-supervised support vector machines, graph-based, and co-training. Semi-supervised learning can be categorized into 5 groups: generative, consistency regularization, graph-based, pseudo-labeling, and hybrid methods.
2.0.2 Semi-supervised semantic segmentation
addresses the challenges of applying semi-supervised learning principles to the task of semantic image segmentation. It is a more realistic application of semi-supervised learning since semantic segmentation involves pixel-level labeling which is a more intensive labeling requirement than classification or detection. The techniques in this field can also be categorized into 5 categories: Adversarial, contrastive learning, consistency regularization, pseudo-labeling, and hybrid methods.
2.0.3 Consistency regularization
The consistency regularization method commonly used in semi-supervised semantic segmentation can be categorized into 4 main groups: Input perturbations, feature perturbations, network perturbations, and the three combined methods. The common goal for all of these sub-groups is to make the model perform consistently across all transformations applied to the unlabelled data.
2.0.4 Mean teacher.
Almost all of the methods in consistency regularization are based on the Mean Teacher model. The Mean Teacher model consists of two main components: the student model and the teacher model. The student model is the main model that is being trained on the available labeled data. Like in standard supervised learning, the student model learns to make predictions based on the labeled examples it sees. For the Teacher model, this is a secondary model whose weights are an exponential moving average (EMA) of the student model’s weights over time. The teacher model is not trained directly with gradient descent. Instead, its weights are updated as the moving average of the student model’s weights after each training step.
3 Methodology
3.1 Problem definition
Semi-supervised semantic segmentation can be formulated as follows: Let $X=\{X_{L},X_{U}\}$ be the dataset, and $X_{L}=\{x_{i}\}_{i=1}^{L}$ be the labeled subset, $Y_{L}=(y_{1},y_{2},\ldots,y_{L})$ is the labels for such set and $X_{U}=\{(x_{i})\}_{i=1}^{U}$ be the unlabeled subset. The assumption is that $L\ll U$. The objective is to minimize the following loss function:
$\min_{\theta}\sum_{x\in X_{L},y\in Y_{L}}\mathcal{L}_{s}(x,y,\theta)+\alpha\sum_{x\in X_{U}}\mathcal{L}_{u}(x,\theta)+\beta\sum_{x\in X}\mathcal{R}(x,\theta)$
$\mathcal{L}_{s}$ denotes the supervised losses with labeled data, $\mathcal{L}_{u}$ is the unsupervised loss and $\mathcal{R}$ is the regularization term. Depending on the technique chosen, this regularization term can be consistency regularization or another designed regularization technique. $\theta$ is the model’s parameters and $\alpha$ and $\beta$ are trade-off terms respectively.
3.2 Self-training (ST)
Similar to Mean teacher, self-training also incorporated the idea of the teacher-student model. But one significant difference is that the teacher and student are trained in two separate stages. The steps for the plain Self-training are as follows:
  1. 1.
    Train the teacher model $T$ on labeled data $X_{L}=\{x_{i}\}_{i=1}^{L}$.
  2. 2.
    Use the model that has just been trained on labeled data to generate predictions on the unlabeled data. These predictions in the context of semantic segmentation can be considered as pseudo-masks. This set of pseudo-masks can be formulated as $\hat{D}_{U}=\left\{(x_{i},T(x_{i}))\right\}_{i=1}^{U}$
  3. 3.
    Extend the labeled data ${D}_{L}=\left\{(x_{i},y_{i})\right\}_{i=1}^{L}$ with the unlabeled ones $\hat{D}_{U}$ by unioning the labeled set with the unlabeled set, then re-train the student model S on this new set.
In [16], the authors propose a stronger baseline for this plain self-training pipeline by incorporating strong data augmentation in the final stage on unlabelled data. The performance upon implementing these is competitive in comparison with other methods at the time.
Algorithm 1 ST Pseudocode
1:Labeled training set $D_{L}=\{(x_{i},y_{i})\}_{i=1}^{L}$, Unlabeled training set $D_{U}=\{u_{i}\}_{i=1}^{U}$
2:Fully trained model $S$
3:Initialize teacher model $T$ with random weights
4:Pre-train $T$ on $D_{L}$ using supervised learning
5:Use model $T$ to infer labels on $D_{U}$ to get $\hat{D}_{U}$
6:Combine $D_{L}$ and $\hat{D}_{U}$ to create an augmented training set $D_{aug}$
7:Re-train student model $S$ on $D_{aug}$
8:return $S$
3.3 Self-training++ (ST++)
Despite the simplicity and impressive results of Self-training, it doesn’t have a mechanism to take into account the pseudo-masks that are extremely incorrect. By re-training on pseudo-masks that are extremely wrong, called unreliable samples, the model’s performance inevitably deteriorates. In contrast, the pseudo-masks that are close to the actual ground truth of the unlabeled image, called reliable samples, play a key role in improving the model’s performance. Therefore, in [16], the authors propose a way to harness the power of reliable images to compensate for unreliable ones. Instead of re-training directly on labeled data and all unlabelled data in the final step, the authors first train the model on labeled data and reliable samples. Then, they use this model to generate pseudo-masks for the unreliable data. Finally, they re-train the model on labeled data, reliable data, and unreliable data.
Same as the strong baseline the authors in [16] for plain self-training, the authors apply strong data augmentation on the unlabeled images whenever re-training on both labeled and unlabeled data.
Algorithm 2 ST++ Pseudocode
1:Labeled training set $D_{L}$, Unlabeled training set $D_{U}$
2:Fully trained student model $S$
3:Train teacher model $T$ on $D_{L}$
4:Obtain pseudo labels for $D_{U}$ using $T$
5:Select a subset of $D_{U}$ with the most reliable pseudo labels
6:Train student model $S$ on $D_{L}$ and the reliable subset of $D_{U}$
7:Re-label the remaining subset of $D_{U}$ using $S$
8:Re-train $S$ on $D_{L}$ and the entire $D_{U}$
9:return $S$
3.4 FixMatch
Inspired by the input perturbation method in the area of consistency-regularization applied in semi-supervised semantic segmentation, FixMatch [17] is a method introduced in 2020 based on the assumption that a model should produce the same prediction for an unlabeled image, regardless of whether it has been weakly or strongly augmented. At the time of its introduction, the FixMatch method outperformed all other methods in datasets such as CIFAR-100, SVHN, and STL-100 [17] in the field of image classification.
3.5 ST++ with FixMatch
Realizing that the ST++ approach has not yet fully utilized consistency regularization approaches, we aspire to incorporate FixMatch among the steps described in the pseudo-code for ST++. Realizing that the first step of training in a fully supervised manner in ST++ can be easily adjusted to incorporate FixMatch. We have done just that.
Algorithm 3 ST++ with FixMatch
1:Labeled dataset $D_{L}=\{(x_{i},y_{i})\}$, Unlabeled dataset $D_{U}=\{u_{j}\}$, Model parameters $\theta$, Confidence threshold $\tau$, Strong augmentation function $S(\cdot)$, Weak augmentation function $W(\cdot)$
2:Initialize: Train model on $L$ for a few epochs
3:for each training epoch do
4:     for each mini-batch do
5:         Draw a batch of labeled samples $\{(x_{i},y_{i})\}$ from $D_{L}$
6:         Apply weak augmentation: $x^{\prime}_{i}=\alpha(x_{i})$
7:         Draw a batch of unlabeled samples $\{u_{j}\}$ from $U$
8:         Apply weak augmentation: $u^{\prime}_{j}=W(u_{j})$
9:         Apply strong augmentation: $u^{\prime\prime}_{j}=S(u_{j})$
10:         Obtain predictions for weakly-augmented unlabeled data: $p_{j}=M(u^{\prime}_{j};\theta)$
11:         Generate pseudo-labels: $\hat{y}_{j}=\arg\max(p_{j})$ if $\max(p_{j})>\tau$ else ignore
12:         Supervised Training Step (incorporating FixMatch):
13:         Train model on $\{(x^{\prime}_{i},y_{i})\}$ using supervised loss
14:         Train model on $\{(u^{\prime\prime}_{j},\hat{y}_{j})\}$ for samples with pseudo-labels, using consistency loss
15:     end for
16:     Update model parameters $\theta$
17:end for
18:return Optimized model parameters $\theta$
4 Experiments
4.1 Dataset
4.1.1 Dataset 1.
[14] Sourced from Jiangxi Tecom Science Corporation in China, comprises 300 images of white blood cells (WBCs), each measuring 120×120 pixels with a color depth of 24 bits. These images were captured using a Motic Moticam Pro 252A camera attached to an N800-D motorized auto-focus microscope. The blood smear samples in this dataset were stained using a novel hematology reagent specifically designed for rapid WBC staining, resulting in images that predominantly exhibit a yellow background.
Figure 1: Dataset1: A sample
4.1.2 Dataset 2
[14] Dataset 2 is composed of 100 color images, each with dimensions of 300×300 pixels, acquired from the CellaVision blog. These images predominantly feature cells in a purple hue and often include numerous red blood cells surrounding the white blood cells. This characteristic sets Dataset 2 apart from Dataset 1.
4.1.3 LISC.
[15] The LISC dataset consists of hematological images taken from the peripheral blood of healthy subjects. The dataset contains images of normal leukocytes, including basophils, eosinophils, lymphocytes, monocytes, and neutrophils, manually segmented and classified by experts. The images, collected from 8 normal subjects and amounting to 400 samples from 100 microscope slides, are in BMP format with a resolution of 720×576 pixels.
4.2 Settings
The labeled set accounts for 1/4 of the size of the training set. This proportion is in accordance with the setting used in the semi-supervised learning setting, such as ST++ [16].
The experiment was conducted on the RTX3090. The experiment was run under two settings: With FixMatch and without FixMatch. Under each of these settings, we employ three frameworks in semi-supervised learning:
  • SupOnly: The model is trained on the labeled dataset only.
  • ST: The model is trained in a standard self-training manner.
  • ST++: The model is trained in a self-traning++ framework according to the pseudo-code 2.
Without consistency, the batch size for Dataset 1, Dataset 2, and LISC was 16. With the consistency incorporated, the batch size for the labeled and the unlabeled data is 16. For Dataset 2 and LISC, the labeled data’s batch size is 16 and the unlabeled data’s batch size is 8. The network architectures are DeepLabV3+, DeepLabv2, and PSPNet. Each network architecture consists of the ResNet-50 and ResNet-101 backbone, but DeepLabv2 only has ResNet-101 backbone.
We will also run two algorithms: Self-Training (ST) and Self-Training++ (ST++)
5 Results
Table 1: The mIOU of the models run under different networks with different backbones.
Model Backbone Backbone Dataset
Dataset 1 Dataset2 LISC
SupOnly ST ST++ SupOnly ST ST++ SupOnly ST ST++
without consistency 89.49 90.71 90.02 86.19 86.43 87.34 75.2 76.67 77.64
DeepLabv3 ResNet-50 with consistency 87.81 90.04 90.69 84.19 87.02 87.37 73.96 76.48 76.49
without consistency 88.86 89.83 89.99 86.03 86.35 87.22 75.95 75.87 76.55
DeepLabv3 ResNet-101 with consistency 86.61 89.8 90.29 85.15 87.28 87.09 75.09 77.02 76.6
without consistency 82.54 84.21 85.01 74.46 79.7 78.74 72.2 75.27 74.84
DeepLabv2 ResNet-101 with consistency 80.08 85.05 85.07 71.47 79.32 80.38 70.47 74.71 75.75
without consistency 80.27 83.08 83.64 75.76 80.2 80.92 71.17 74.69 75.12
PSPNet ResNet-50 with consistency 79.02 83.44 83.76 74.95 80.31 81.79 71.53 75.5 75.39
without consistency 81.56 83.93 84.3 77.76 80.39 81.35 72.91 75.19 75.17
PSPNet ResNet-101 with consistency 80.69 84.56 84.58 77.33 79.8 80.11 71.35 74.96 74.82
From the above results 1, there can be two conclusions:
  • There is a performance improvement when transitioning from Supervised-Only (SupOnly) to Self-Training (ST) and further to ST++, with ST++ demonstrating the highest level of performance.
  • With FixMatch incorporated, the performance on ST and ST++ framework improves in certain scenarios.
6 Discussion
6.1 Performance on the supervised training stage when incorporating FixMatch
When incorporating FixMatch into the supervised training stage, there was a noticeable decrease in performance on the labeled images. This is a commonly seen trade-off when we try to do two tasks in 1 stage.
Table 2: Noticeable decrease when incorporating FixMatch into the model.
Dataset1 Dataset2 LISC
$\times$ $\times$ $\times$
DeepLabv3+ - ResNet-50 89.49 87.81 86.19 84.19 75.2 73.96
DeepLabv3+ - ResNet-101 88.86 86.61 86.03 85.15 75.95 75.09
DeepLabv2 - ResNet101 82.54 80.08 74.46 71.47 72.2 70.47
PSPNet - ResNet50 80.27 79.02 75.76 74.95 71.17 71.53
PSPNet - ResNet101 81.56 80.69 77.76 77.33 72.91 71.35
However, the performance in the later stage remains the same or even improves, suggesting that the trade-off in the first stage is sometimes necessary to improve the performance when extending the labeled images with unlabeled ones.
6.2 The quality of the pseudo-masks produced when incorporating FixMatch
The quality of the pseudo-masks between the two settings of FixMatch and non-FixMatch approaches varies. On the one hand, the pseudo-masks of easy, reliable images remain essentially the same between the two settings.
Figure 2: Sample of Ground truth mask, Non-FixMatch, and FixMatch mask generated from the LISC dataset.
However, for more complicated, unreliable images, the masks generated can have a large deviation from the ground truth mask.
Figure 3: Deviations from the Ground truth mask of Non-FixMatch and FixMatch mask generated from the LISC dataset.
This suggests that incorporating FixMatch into the model has yet to be able to take on the nuances presented in challenging images.